The equation for the combustion of butane, C4H10, is:
2C4H10 (g)+13O2 (g)→8CO2 (g)+10H2O(g)
Which one of the following situations generates the smallest amount of heat? Motivate your answer.
a) Burning one mole of butane.
b) Reacting one mole of oxygen with excess butane.
c) Producing one mole of carbon dioxide by burning butane.
d) Producing one mole of water by burning butane.
1
Expert's answer
2020-04-06T11:59:53-0400

The more butane is burnt, the more heat is produced. So calculating what quantites of butane undergo reaction in each case will show what amounts of heat are produced.

1) n(C4H10) = 1 mol

2) n(C4H10) = "{\frac {2} {13}}" mol accroding to equation, as n(O2) = 1 mol and n(C4H10) = "{\frac {2} {13}}"*n(O2)

3) n(C4H10) = "{\frac {1} {4}}" mol according to equation, as n(CO2 produced) = 1 mol and n(C4H10) = "{\frac {2} {8}}"*n(CO2)

4) n(C4H10) = "{\frac {1} {5}}" mol according to equation, as n(H2O produced) = 1 mol and n(C4H10) = "{\frac {2} {10}}"*n(H2O)

the least amount of butane involved is in the second case, so the least amount of heat generated in 2 situation